OpenStudy (anonymous):

lim x->1 ln(x)/(sin(pi*x)

OpenStudy (anonymous):

l'Hopital's?

OpenStudy (anonymous):

\[\frac{ \frac{ 1 }{ x } }{\pi \cos \pi x }\]